OpenStudy (anonymous):

Suppose a special-effects team made a model tower for a movie. The height of the model was 9.2 meters and its width was 6 meters. What is the height of the actual tower if its width is 996 meters? Round your answer to the nearest meter.

OpenStudy (valpey):

There was no actual tower. It was just a movie. Who would build a tower that was 996 meters wide and 996*9.2/6 meters tall?

OpenStudy (lgbasallote):

use ratio and proportion here. you're looking for the height so let that be equal to x 9.2 : 6 = x : 996 \[\frac{9.2}{6} = \frac{x}{996}\] \[\frac{9.2}{6} \times 996 = x\]